### v1

@@ VI-A Evaluation setup
We conduct a comparative study in which, first, we drive MTs by an unadvised RL agent, and then, we provide advice at various levels of uncertainty. We use the $12\times 12$ scaled-up, randomized version of the Frozen Lake environment (Fig. 1). The map contains 20% Hole tiles, which is a standard ratio in OpenAI Gym’s Frozen Lake implementation. The exact map is visualized in the replication package. We assess the performance of RL agents by the cumulative reward they collect. We measure both the amount of the cumulative reward and the pace at which the RL agent collects it. Our hypothesis is that human guidance has measurable positive impact in both aspects, even at moderate levels of uncertainty.
@@
The RL-specific parameters of our experiments are shown Tab. I. For the RL method, we choose discrete policy gradient [sutton2018reinforcement] as policy-based methods explicitly represent the policy and allow for the direct investigation of the effects of advice. Learning rate $\alpha$ affects how much the agent updates its policy parameters during training. The closer it is to 00, the smaller the policy parameter updates, and the closer it is to $1$, the larger the policy parameter updates. Discount factor $\gamma$ determines how much the agent values future rewards. The closer $\gamma$ is to 00, the more the agent prioritizes immediate rewards, and the closer it is to $1$, the more the agent values long-term rewards. Each experiment runs for a number of episodes. An episode starts by resetting the environment and placing the agent in the initial position. Within the same experiment, the policy is continuously maintained. An episode lasts until the agent either finds a terminating state (hole or goal) or reaches the number of maximum steps per episode. To mitigate threats to validity, we tuned these hyperparameters to favor the learning dynamics of the unadvised agent.
@@
Source of advice {Oracle, Single human}
@@
We guide advised agents by different sources of advice and advice quotas. To form a ground truth, we first experiment with an idealized oracle with full information of the problem space, that provides advice based on pre-programmed rules. The oracle has two quotas; one where it provides advice about all states (100%) and another where it provides advice about the holes and goal on the map (20%).
@@
Note about additional experiments in the replication package
Our replication package contains data of additional experiments, in which two cooperating humans guide the RL agent. Due to space restrictions, we do not present those results here, but we note that agents advised by cooperating humans outperform the unadvised agent, and cooperating humans can obtain similar performance to a single human advisor.
@@ VI-B Results and key takeaways
We observe that the oracle with 100% quota has the best overall performance, accumulating the most cumulative reward for all but one $u=0.0$, where the oracle with 20% quota is slightly better. The oracles with 100% and 20% quota, and the human advisor with 10% have similar performance for low uncertainty $u=0.0$ and $u=0.2$. This is indicated by the cumulative reward, and the similar slope in the log plots. While the oracle with 100% advice quota (Fig. 5(a)) significantly outperforms the human advisor with 10% quota (Fig. 5(c)) at low levels of uncertainty ($u\leq 0.4$), this difference is not significant at higher levels of uncertainty ($u\geq 0.6$). That is, in uncertain situations, a human advisor, even with lower advice quota does not perform systematically worse than a hypothetical oracle. This finding highlights the utility of human advice in uncertain situations. This trend is even more pronounced when the oracle is given less advice quota (Fig. 5(b)). With 20% quota, the oracle does not significantly outperform a human advisor with 10% quota when $u>0.0$.
At low levels of uncertainty ($u\leq 0.4$) a human advisor with 10% quota significantly outperforms the human advisor with 5% quota. However, at higher levels of uncertainty ($u\geq 0.6$), the difference is not significant. These findings suggest that different uncertainty levels require different advice strategy from human advisors. Specifically, in rather certain situations, more advice is justified; while in rather uncertain situations, advisors should be more frugal with their advice.
@@
[…]
innertopmargin=4pt, frametitleaboveskip=-frametitlealignment= {mdframed} Our observations indicate that human guidance contributes to significant performance improvement in RL-based inference of complex MTs, even at moderate-to-high uncertainty. Human advisors, even with substantially lower advice quotas, can perform at the level of an oracle, and might even outperform it. This results in faster and more efficient inference of complex MTs.
@@ VI-C1 Assumptions and limitations
[…]
@@ VI-C2 Threats to validity
A fortunately aligned reward structure may result in better, albeit non-systemic performance of the RL agent, threatening the internal validity of our study. To mitigate this threat, (i) we have chosen the least supportive reward function in which the goal reward is minimal, there are no intermediate rewards, and terminating states do not provide negative feedback; and (ii) tuned the hyperparameters to favor the learning dynamics of the unadvised agent. To assess the external validity of our study, we remark that we only tested policy-based RL. While it is reasonable to expect the method to be applicable to other RL methods (e.g., value-based RL, a popular choice in MDE [barriga2022parmorel, eisenberg2021towards, sharbaf2022automatic]), our study does not provide evidence for this. Similarly, a minor threat to external validity stems from the technological choices, especially the MT platform. We rely on the reactive capabilities of Viatra, and it is reasonable to expect that MT frameworks with support for similar MT execution semantics can adopt our approach.
@@ VII-B1 Model synchronization and inconsistency management
These problems are related and both often require identifying complex MTs [mens2007incremental, elhamlaoui2018mde, david2016engineering]. In these settings, the RL state encapsulates morphisms over the union of the source and target model, and the reward is typically implied by the negative distance between the source and target models. The lower the distance, the better the solution. Termination semantics are asserted to states that minimize the distance or, equivalently, maximize the reward. Distance can be measured at the syntactic level (e.g., by counting inconsistent model elements) [marchezan2024tool] or at the semantic level (e.g., by counting inconsistent properties that are not being jointly satisfied by the two models) [david2023model].
@@
Actions are typically expressed as domain-specific MTs, but are limited in complexity. Our approach allows for inferring complex MTs that allow for more sophisticated synchronization semantics between models. Human guidance is of particular value in these problems as multi-domain settings challenge the understanding of how exactly consistency should be restored [jongeling2023uncertainty, dagenais2024driving].
@@ VII-B2 Model repair and refactoring
This class of problems is similar to model synchronization, however, there are some key differences. RL state in this case encapsulates only a single model’s morphisms, and reward cannot be automatically calculated due to the lack of a suitable oracle. That is, it cannot be structurally defined when a refactoring is successful or sufficient. Instead, the reward structure must be defined by the domain expert. Typically, qualitative metrics are used for this purpose, either defined at the general syntactic level (e.g., number of dangling edges or isolated components) [taentzer2017change-preserving] or at the domain-specific syntactic level [syriani2019domain-specific]. Occasionally, personalized metrics are considered [barriga2022parmorel].
@@
The challenge is similar to the previous class of problems: actions are usually expressed as domain-specific MTs with limited complexity. Here again, an approach such as ours would allow for inferring complex MTs that, in turn, enable more complex refactorings and model repair actions. Human guidance helps prevent models to grow out of size under the premise of repair.
@@ VII-B3 Design space exploration
Design space exploration (DSE) is the process of identifying favorable candidate designs under various constraints and soft objectives [herzig2017model-transformation-based]. Starting from an initial design, a DSE process gradually applies MTs to uncover alternative designs. In such scenarios, the state encapsulates a specific graph morphism in a model instance. Actions are typically defined in domain-specific terms [hegedus2015model] rather than domain-agnostic CRUD primitives. The reward structure is given through constraints and soft objectives.
@@
Optimal solutions are often found deep in the design space and are computationally demanding to identify. Our approach can aid the computational load of DSE by learning complex MTs that have led to beneficial design alternatives. These complex MTs, then, can be used as fixed points in the design space from which subsequent DSE executions may continue deeper explorations. Human guidance helps by providing known favorable design decisions and by that, drive the RL agent to favorable states and accelerate the inference of complex MTs. The challenge of uncertainty in DSE has been recognized before [foldiak2024probabilistic]. Our approach allows for the explicit modeling and management of uncertainty that stems from human guidance.
@@ VII-C1 Domain-specific languages for guidance
A key advantage of opinion-based guidance is the rapid pace at which opinions emerge. To leverage this advantage to its fullest, intuitive languages are needed to express opinions. Domain-specific languages (DLSs) [schmidt2006model] have been widely used to capture expert opinions at high levels of abstraction, narrowing the gap between the modeling language and human cognition. However, DSLs typically operate with reduced syntax to remain truly specific to the problem domain, and therefore, automated methods for DSL engineering are in high demand [bucchiarone2020grand]. This is true for guided RL as well. We envision language workbenches that allow for the automated construction of DSLs through which advice can be provided by domain concepts. In the running example (Sec. III), one might want to use descriptive quantifiers, such as “hazardous” or “beneficial” instead of relying on an integer scale. Unfortunately, there is little research on domain-specific languages for RL, and existing languages mostly focus on programming RL procedures [molderez2019marlon] rather than expressing opinions at high cognitive levels. We see plenty of room for research targeting domain-specific modeling languages for RL, inspired by the advancements in the model-driven engineering community [kulagin2022ontology, wu2005automated].
@@
Our approach can be aided by domain-specific tool support at key points. For example, the domain expert needs a domain-specific editor with an intuitive user interface to efficiently provide advice. We implemented the prototype of such an editor on top of Eclipse and Xtext, as discussed in Sec. V-A. Instead of case-by-case tool development, syntax-directed editors could be generated from high-level specifications of the RL problem and the application domain [syriani2021generation], and equipped with automatically generated specialized concrete syntax [benchaaben2024intelligent]. Adaptive modeling languages [delara2025adaptive] offer an apt solution through a product line-inspired approach to customizing DSLs to different problems.
Additional tooling considerations relate to model management and RL execution. In our work, we executed the RL agent’s learning process on an instance model that should likely be prevented from automated complex MTs until the RL agent is trained. Instead, an improved version of our tool could create replicas of the real model and treat those as sandbox models for the agent to train. The result of the training, i.e., the policy, then, can be applied on the real model. Such utility functionality will pave the way for adoption.
@@ VII-C3 Interactive advice
In our experiments, we assumed that advice is provided once, before the agent begins exploring the environment. However, given adequate interaction protocols, our approach can accommodate advice at any point during the RL process. We see opportunities in developing more interactive advice protocols through advanced human-computer interactions (HCI) [shajari2025bridging]. MDE can be of high value in such situations, as it can support domain-specific interfaces that integrate user models in the system design. This would allow for interactive advising systems that are tailored to the user’s level of expertise about the problem space [abrahao2017user]. Blended modeling—the activity of interacting seamlessly with a single model (i.e., abstract syntax) through multiple notations (i.e., concrete syntaxes), allowing a certain degree of temporary inconsistencies [david2023blended]—can be of high utility in facilitating seamless advising workflows in multi-view settings.
Finally, interaction protocols amenable to verification and validation can be generated, e.g., from state chart models [harel1987statecharts] or sequence diagrams [kundu2013automatic].
@@ VIII Conclusion
We presented a method for engineering complex model transformation (MT) sequences by reinforcement learning (RL), guided by potentially uncertain human advice. Our results show that human advice, even if uncertain, can be a major contributor to more efficient MT development. Our results open the door for the RL-based human-in-the-loop treatment of key MDE problems, e.g., model synchronization and design-space exploration. We envision MT platform developers to adopt techniques such as ours, and promote RL-backed MTs to first-class citizens in their platforms. This will enable efficient MT development, execution, and maintenance.

### v2

@@ VI-A Evaluation setup
We conduct a comparative study in which, first, we drive MTs by an unadvised RL agent, and then, we provide advice at various levels of uncertainty. We use the $12\times 12$ scaled-up, randomized version of the Frozen Lake environment (Fig. 1). The map contains 20% Hole tiles, which is a standard ratio in OpenAI Gym’s Frozen Lake implementation. The exact map is visualized in the replication package. We assess the performance of RL agents by the cumulative reward they collect. We measure both the amount of the cumulative reward and the pace at which the RL agent collects it. Our hypothesis is that human guidance has a measurable positive impact on both aspects, even at moderate levels of uncertainty.
@@
The RL-specific parameters of our experiments are shown in Tab. I. For the RL method, we choose discrete policy gradient [4] as policy-based methods explicitly represent the policy and allow for the direct investigation of the effects of advice. Learning rate $\alpha$ affects how much the agent updates its policy parameters during training. The closer it is to 0, the smaller the policy parameter updates, and the closer it is to $1$, the larger the policy parameter updates. Discount factor $\gamma$ determines how much the agent values future rewards. The closer $\gamma$ is to 0, the more the agent prioritizes immediate rewards, and the closer it is to $1$, the more the agent values long-term rewards. Each experiment runs for a number of episodes. An episode starts by resetting the environment and placing the agent in the initial position. Within the same experiment, the policy is continuously maintained. An episode lasts until the agent either finds a terminating state (hole or goal) or reaches the number of maximum steps per episode. To mitigate threats to validity, we tuned these hyperparameters to favor the learning dynamics of the unadvised agent.
@@
Source of advice
{Oracle, Single human}
@@
We guide advised agents by different sources of advice and advice quotas. To form a ground truth, we first experiment with an idealized oracle with full information of the problem space, which provides advice based on pre-programmed rules. The oracle has two quotas; one where it provides advice about all states (100%) and another where it provides advice about the holes and goal on the map (20%).
@@
[…]
[width=]figures/results/plot_allRewardData_linear.pdf
@@
[…]
[width=]figures/results/plot_holesAndGoalRewardData_linear.pdf
@@
[…]
[width=]figures/results/plot_human10RewardData_linear.pdf
@@
[…]
[width=]figures/results/plot_human5RewardData_linear.pdf
@@
[…]
[width=]figures/results/plot_human10RewardData_log.pdf
@@
[…]
[width=]figures/results/plot_human5RewardData_log.pdf
@@
Note about additional experiments
The replication package contains additional data from experiments with two cooperating humans guiding the RL agent. Due to space restrictions, those results are not presented, but we note that agents advised by cooperating humans outperform the unadvised agent and cooperating humans with lower quotas obtain similar performance to a single human advisor.
@@ VI-B Results and key takeaways
We observe that the oracle with 100% quota has the best overall performance, accumulating the most cumulative reward for all but one $u=0.0$, where the oracle with 20% quota is slightly better. The oracles with 100% and 20% quota, and the human advisor with 10% have similar performance for low uncertainty $u=0.0$ and $u=0.2$. This is indicated by the cumulative reward and the similar slope in the log plots. While the oracle with 100% advice quota (Fig. 5(a)) significantly outperforms the human advisor with 10% quota (Fig. 5(c)) at low levels of uncertainty ($u\leq 0.4$), this difference is not significant at higher levels of uncertainty ($u\geq 0.6$). That is, in uncertain situations, a human advisor, even with a lower advice quota, does not perform systematically worse than a hypothetical oracle. This finding highlights the utility of human advice in uncertain situations. This trend is even more pronounced when the oracle is given less advice quota (Fig. 5(b)). With 20% quota, the oracle does not significantly outperform a human advisor with 10% quota when $u>0.0$.
At low levels of uncertainty ($u\leq 0.4$) a human advisor with 10% quota significantly outperforms the human advisor with 5% quota. However, at higher levels of uncertainty ($u\geq 0.6$), the difference is not significant. These findings suggest that different uncertainty levels require different advice strategies from human advisors. Specifically, when advisors are certain, more advice is justified; however, when uncertain, advisors should be more frugal with their advice.
@@
Our observations indicate that human guidance contributes to significant performance improvement in RL-based inference of complex MTs, even at moderate-to-high uncertainty. Human advisors, even with substantially lower advice quotas, can perform at the level of an oracle, and might even outperform it. This results in faster and more efficient inference of complex MTs.
@@ VI-C1 Assumptions and limitations
We implicitly assumed that the reward structure could be effectively formulated. In some MDE applications, this is true. For example, [9] formulate reward by a user-preferred quality characteristic, such as maintainability and understandability. Still, formulating reward might be challenging in open problems, such as design space exploration, in which reward often must be asserted to hypothetical, yet-to-encounter model configurations.
@@ VI-C2 Threats to validity
A fortunately aligned reward structure may result in better, albeit non-systemic performance of the RL agent, threatening the internal validity of our study. To mitigate this threat, (i) we have chosen the least supportive reward function in which the goal reward is minimal, there are no intermediate rewards, and terminating states do not provide negative feedback; and (ii) tuned the hyperparameters to favor the learning dynamics of the unadvised agent. To assess the external validity of our study, we remark that we only tested policy-based RL. While it is reasonable to expect the method to be applicable to other RL methods (e.g., value-based RL, a popular choice in MDE [9, 6, 35]), our study does not provide evidence for this. Similarly, a minor threat to external validity stems from the technological choices, especially the MT platform. We rely on the reactive capabilities of Viatra, and it is reasonable to expect that MT frameworks with support for similar MT execution semantics can adopt our approach.
@@ VII-B1 Model synchronization and inconsistency management
These problems are related and both often require identifying complex MTs [36, 37, 38]. In these settings, the RL state encapsulates morphisms over the union of the source and target models, and the reward is typically implied by the negative distance between the source and target models. The lower the distance, the better the solution. Termination semantics are asserted to states that minimize the distance or, equivalently, maximize the reward. Distance can be measured at the syntactic level (e.g., by counting inconsistent model elements) [39] or at the semantic level (e.g., by counting inconsistent properties that are not being jointly satisfied by the two models) [40].
@@
Actions are typically expressed as domain-specific MTs, but are limited in complexity. Our approach allows for inferring complex MTs that allow for more sophisticated synchronization semantics between models. Human guidance is of particular value in these problems as multi-domain settings challenge the understanding of how exactly consistency should be restored [24, 41].
@@ VII-B2 Model repair and refactoring
This class of problems is similar to model synchronization; however, there are some key differences. RL state in this case encapsulates only a single model’s morphisms, and reward cannot be automatically calculated due to the lack of a suitable oracle. That is, it cannot be structurally defined when a refactoring is successful or sufficient. Instead, the reward structure must be defined by the domain expert. Typically, qualitative metrics are used for this purpose, either defined at the general syntactic level (e.g., number of dangling edges or isolated components) [42] or at the domain-specific syntactic level [30]. Occasionally, personalized metrics are considered [9].
@@
The challenge is similar to the previous class of problems: actions are usually expressed as domain-specific MTs with limited complexity. Here again, an approach such as ours would allow for inferring complex MTs that, in turn, enable more complex refactorings and model repair actions. Human guidance helps prevent models from growing out of size under the premise of repair.
@@ VII-B3 Design space exploration
Design space exploration (DSE) is the process of identifying favorable candidate designs under various constraints and soft objectives [43]. Starting from an initial design, a DSE process gradually applies MTs to uncover alternative designs. In such scenarios, the state encapsulates a specific graph morphism in a model instance. Actions are typically defined in domain-specific terms [10] rather than domain-agnostic CRUD primitives. The reward structure is given through constraints and soft objectives.
@@
Optimal solutions are often found deep in the design space and are computationally demanding to identify. Our approach can aid the computational load of DSE by learning complex MTs that have led to beneficial design alternatives. These complex MTs, then, can be used as fixed points in the design space from which subsequent DSE executions may continue deeper explorations. Human guidance helps by providing known favorable design decisions and by that, drive the RL agent to favorable states and accelerate the inference of complex MTs. The challenge of uncertainty in DSE has been recognized before [44]. Our approach allows for the explicit modeling and management of uncertainty that stems from human guidance.
@@ VII-C1 Domain-specific languages for guidance
A key advantage of opinion-based guidance is the rapid pace at which opinions emerge. To leverage this advantage to its fullest, intuitive languages are needed to express opinions. Domain-specific languages (DLSs) [33] have been widely used to capture expert opinions at high levels of abstraction, narrowing the gap between the modeling language and human cognition. However, DSLs typically operate with reduced syntax to remain truly specific to the problem domain, and therefore, automated methods for DSL engineering are in high demand [45]. This is true for guided RL as well. We envision language workbenches that allow for the automated construction of DSLs through which advice can be provided by domain concepts. In the running example (Sec. III), one might want to use descriptive quantifiers, such as “hazardous” or “beneficial” instead of relying on an integer scale. Unfortunately, there is little research on domain-specific languages for RL, and existing languages mostly focus on programming RL procedures [46] rather than expressing opinions at high cognitive levels. We see plenty of room for research targeting domain-specific modeling languages for RL, inspired by the advancements in the model-driven engineering community [47, 48].
@@
Our approach can be aided by domain-specific tool support at key points. For example, the domain expert needs a domain-specific editor with an intuitive user interface to efficiently provide advice. We implemented the prototype of such an editor on top of Eclipse and Xtext, as discussed in Sec. V-A. Instead of case-by-case tool development, syntax-directed editors could be generated from high-level specifications of the RL problem and the application domain [49], and equipped with automatically generated specialized concrete syntax [50]. Adaptive modeling languages [51] offer an apt solution through a product line-inspired approach to customizing DSLs to different problems.
Additional tooling considerations relate to model management and RL execution. In our work, we executed the RL agent’s learning process on an instance model that should likely be prevented from automated complex MTs until the RL agent is trained. Instead, an improved version of our tool could create replicas of the real model and treat those as sandbox models for the agent to train. The result of the training, i.e., the policy, can be applied on the real model. Such utility functionality will pave the way for adoption.
@@ VII-C3 Interactive advice
In our experiments, we assumed that advice is provided once, before the agent begins exploring the environment. However, given adequate interaction protocols, our approach can accommodate advice at any point during the RL process. We see opportunities in developing more interactive advice protocols through advanced human-computer interactions (HCI) [52]. MDE can be of high value in such situations, as it can support domain-specific interfaces that integrate user models in the system design. This would allow for interactive advising systems that are tailored to the user’s level of expertise about the problem space [53]. Blended modeling—the activity of interacting seamlessly with a single model (i.e., abstract syntax) through multiple notations (i.e., concrete syntaxes), allowing a certain degree of temporary inconsistencies [54]—can be of high utility in facilitating seamless advising workflows in multi-view settings.
Finally, interaction protocols amenable to verification and validation can be generated, e.g., from state chart models [55] or sequence diagrams [56].
@@ VIII Conclusion
We presented a method for engineering complex model transformation (MT) sequences by reinforcement learning (RL), guided by potentially uncertain human advice. Our results show that human advice, even if uncertain, can be a major contributor to more efficient MT development. Our results open the door for the RL-based human-in-the-loop treatment of key MDE problems, e.g., model synchronization and design-space exploration. We envision MT platform developers adopting techniques such as ours to promote RL-backed MTs to first-class citizens in MT platforms. This will enable efficient MT development, execution, and maintenance.